OpenStudy (anonymous):

rationalize the denominator and simplify 2+√3/5-√3

OpenStudy (anonymous):

\[\frac{2+\sqrt{3}}{5-\sqrt{3}}\] \[=\frac{2+\sqrt{3}}{5-\sqrt{3}}\times \frac{5+\sqrt{3}}{5+\sqrt{3}}\] \[=\frac{13+7\sqrt{3}}{22}\]